# ESC-Eval: Evaluating Emotion Support Conversations in Large Language Models

### v1

@@ Abstract
Emotion Support Conversation (ESC) is a crucial application, which aims to reduce human stress, offer emotional guidance, and ultimately enhance human mental and physical well-being. With the advancement of Large Language Models (LLMs), many researchers have employed LLMs as the ESC models. However, the evaluation of these LLM-based ESCs remains uncertain. Inspired by the awesome development of role-playing agents, we propose an ESC Evaluation framework (i.e., ESC-Eval), which uses a role-playing agent to interact with ESC models, followed by a manual evaluation of the interactive dialogues. In detail, we first re-organize 2,801 role-playing cards from seven existing datasets to define the roles of the role-playing agent. Second, we train a specific role-playing model called ESC-Role which behaves more like a confused person than GPT-4. Third, through ESC-Role and organized role cards, we systematically conduct experiments using 14 LLMs as the ESC models, including general AI-assistant LLMs (e.g., ChatGPT) and ESC-oriented LLMs (e.g., ExTES-Llama). We conduct comprehensive human annotations on interactive multi-turn dialogues of different ESC models. The results show that ESC-oriented LLMs exhibit superior ESC abilities compared to general AI-assistant LLMs, but there is still a gap behind human performance. Moreover, to automate the scoring process for future ESC models, we developed ESC-RANK, which trained on the annotated data, achieving a scoring performance surpassing 35 points of GPT-4. Our data and code are available at https://github.com/haidequanbu/ESC-Eval.
@@
Haiquan Zhaothanks: Work done during internship at Shanghai Artificial Intellignece Laboratory1,2   Lingyu Li2   Shisong Chen1   Shuqi Kong2    Jiaan Wang 1 Kexing Huang2   Tianle Gu2   Yixu Wang2   Jian Wang2   Dandan Liang2 Zhixu Lithanks: Correspondence to: Zhixu Li<zhixuli@fudan.edu.cn> and Yan Teng <tengyan@pjlab.org.cn>1  Yan Teng22footnotemark: 22    Yanghua Xiao1   Yingchun Wang2 1 Fudan University 2 Shanghai Artificial Intelligence Laboratory
@@ 2.2 Role Card Acquisition
To ensure the diversity of character cards, drawing inspiration from ESConvLiu et al. (2021), ExTESZheng et al. (2023b), and the Life Events ScaleWethington (2016), we first constructed a classification system consisting of three hierarchical layers and encompassing 37 categories. Then we propose reconstructing role cards in open-resources data and identifying each role card within each category. The construction of this procedure involves three primary steps. First, we collect 7 open-source datasets that cover a wide range of potential user roles. Then we utilize GPT-4 to extract roles from these datasets and filter out low-quality role cards, followed by human filtering. Finally, we employ a manual annotation process to ensure the quality of the role cards and classify them into their respective tertiary categories. We will introduce each step in the following, and more details can be found in Appendix A.
@@ 2.2.1 Dataset collection
To obtain a diverse set of character cards, we conducted a comprehensive investigation into existing datasets in the field of emotion support and mental health datasets. Subsequently, we selected seven datasets as the source datasets for this study. The open-source datasets utilized are listed in Appendix A.
@@ 2.2.2 User cards extraction and filtering
After obtaining these datasets, we encountered both Multi-turn Dialogue (MD) datasets and single-turn Question-and-Answer (QA) datasets. To extract user profiles from these diverse datasets, we employed different prompts for QA and MD datasets using GPT-4 for the initial extraction. The utilization of GPT-4 in this process incurred approximately a cost of \$120. After acquiring the initial character cards, we employ GPT-4 to conduct an initial filtration process on role cards, eliminating those that solely consist of emotions without any associated events. The utilization of GPT-4 in this process incurred an approximate cost of \$70. The prompt used in this section can be found in Appendix A. After the filtering process of GPT-4, we apply a human filter to ensure the quality of these cards.
@@ 2.2.3 Manual annotation and correction
After obtaining role cards that had undergone preliminary screening by GPT-4 and human filter, we employed a two-stage approach involving crowdsourced annotation followed by manual correction to ensure the quality of the role cards.
Crowd workers annotation In the above section, we developed a comprehensive three-tier classification system comprising a total of 37 categories of real-life questions, which are listed in Table 9. Based on this classification, the crowd workers are instructed to annotate the filtered character cards with their corresponding tertiary classifications. This requirement is motivated by the need for convenient evaluation and the quality of role-playing. Additionally, we request the crowd workers to label the high-quality and medium-quality character cards within the dataset. The annotation rules and classification system for annotation can be found in Appendix A. The estimated duration for the annotation phase required is approximately 10 days with 10 crowdsourcing workers.
Human correction Upon completion of the first-stage crowdsourced annotation, we proceeded with a second-stage manual correction. We requested persons (the authors of this paper) who are more familiar with this task to conduct an examination of the annotations for each role card, rectifying any incorrect categorizations and addressing issues pertaining to the quality of the role cards
Following the two-stage process of crowdsourced annotation and manual correction, the role cards representing various real-world individuals with different problems were successfully reconstructed. The data analysis of these role cards is listed in Appendix A.
@@ 2.3 ESC-Role
To construct a more robust role-playing model, we trained a specific role-playing agent called ESC-Role using both general data and data specific to ESC scenarios for ESC-Eval. The following sections outline the steps involved in training and evaluating this model.
@@ 2.3.1 Data Collection
Using the same procedure as in Section 2.2, we selected Smile, ESConv, and ExTES datasets mentioned previously to collect ESC scenarios data. We employed methods including extraction through GPT-4, filtering with GPT-4, and manual filtering to extract role cards from multi-turn dialogues. Resulting in a total of 3,390 role-playing data which consist of a role card and a corresponding dialogue. The role cards were used as system prompts for model training. To further enhance the model’s robust role-playing ability, we filtered five role-playing datasets consisting of multi-turn dialogues from Huggingface [^1] https://huggingface.co/. After processing, we acquired 14K role-playing data instances, consisting of both general role-playing instruction data and ESC role-playing data.
@@ 2.3.2 Implementation and Evaluation Metric
Due to the inclusion of both English and Chinese in the character cards, we selected Qwen1.5-14B-Chat as our base model. We adopted LoRAHu et al. (2021) parameter-efficient Finetuning on the dataset collected above. We compared ESC-Role with some state-of-the-art role-playing agents like GPT-4 and BaichuanNPC, these agents are API-based LLMs, and we conducted all kinds of prompts like Chain-of-Thought (CoT)Wei et al. (2022) and In-Context-Learning (ICL)Min et al. (2022), more details can refer to Appendix B.
@@ 2.3.3 Evalution Results
In addition, we selected pairs that had the best performance with different role-playing models and the source multi-turn dialogue data of role cards. We manually evaluated which dialogue more closely resembled real human conversations. The results are shown in Figure 3. From the figure, it can be observed that it is difficult for humans to distinguish between the results generated by ESC-Role and the results from the original data. Both of them outperformed GPT-4 and Baichuan-NPC, demonstrating the effectiveness of using ESC-Role for role-playing in ESC-Eval.
@@ 3 Evaluation
In this section, we conducted evaluations on 14 general LLMs and domain-specific LLMs on ESC-Eval. We first introduce the models for evaluation. Then we display our experimental results. Finally, we display the details of our scoring model ESC-RANK.
@@ 3.1 Evaluating models
We have selected 14 models for evaluation, including closed-source, open-source, and domain-specific models, which are as follows:
@@
To facilitate a more accurate comparison of the capabilities of various models, we have chosen models of similar magnitudes, such as the 6B/7B/8B model parameter sizes for comparison.
@@ 3.2 Evaluation Results
Based on pre-defined dimensions, we conducted a comprehensive manual assessment, and the results are presented in Table 2. Both in English and Chinese ESC conditions, domain-specific LLMs (ChatCounselor and EmoLLM), respectively achieved the best results. From Table 2 above, in the comparison between general models and domain-specific models, the general models perform better in terms of fluency, expression diversity, and emotional comfort skills. This can be attributed to their highly structured output, such as phrases like “I understand you very well, it is very normal to feel …, here are some possible suggestions: ” The general models generate a large amount of text, scoring high in terms of advice effectiveness and expression diversity. Besides, due to Larger scale parameters, the API-based models exhibit greater knowledge of emotional comfort, with GPT-4 and ChatGPT demonstrating the highest proficiency. However, these models perform poorly in terms of human-like and human-centric responses, as users in this context expect replies that are more humanized and possess greater human-like qualities. In the comparison of domain-specific models, MindChat, SoulChat, and EmoLLM, which were not fine-tuned in English, showed inferior fluency. On the other hand, ExTES-llama and ChatCounselor performed well. ExTES was fine-tuned with data generated by ChatGPT, while ChatCounselor was fine-tuned using real psychological counseling data, exhibiting superior performance. From Table 2 bottom, the general models perform well in terms of expressing diversity and providing effective suggestions. Trained on diverse and abundant data, EmoLLM exhibits excellent performance across multiple dimensions among the various domain-specific models. Other domain-specific models, due to their remarkable human-like qualities and human convenience, surpass the general models. However, there is still room for improvement in terms of emotional support knowledge, and significant potential exists for enhancing human convenience. It is worth noting that MindChat, trained on bilingual data, not only demonstrates strong Chinese language proficiency but also exhibits commendable English language capabilities.
@@ 3.3 Correlation Analysis
To validate the effectiveness of ESC-Eval, we randomly selected 20 instances from the ESConv dataset. We chose three categories from the target model and included five different models for correlation analysis. These models were subjected to interactions with human evaluators who model seekers looking for help. And they are asked to provide ratings upon completion of the interactions, according to human evaluation methods in other papers. The human-rated scores were considered as the optimal evaluation method, and we conducted a correlation analysis between various automatic evaluation methods and the ESC-Eval method. The results are presented in Table 3.
@@
From Table 3, it can be observed that ESC-Eval exhibits the best correlation with the evaluation metrics, except for the Fluency and Empathy indicators. In terms of Fluency, automated metrics outperform ESC-Eval, we analyze that during manual annotation, human annotators may exhibit some bias towards the fluency of segmented statements generated by a general model which significantly deviates from the ESConv dataset. It has been observed that human annotators tend to prefer naturally expressed content, leading to relatively lower manual ratings for the fluency outputs of general models. At the same time, the content generated by the general model is quite different from that of ESconv, and the automation metric is also very low. As a result, there is a strong correlation between automated evaluation metrics and humans. However, in ESC-Eval all models perform well on fluency due to the capability of LLMs, leading to low correlation. A similar phenomenon is observed for the Empathy indicator, where although there is some correlation, it is due to the alignment process that most LLMs undergo, which enables them to display decent comforting abilities and analytical skills. In terms of the overall average metric, ESC-Eval demonstrates the most significant correlation compared to the automated metrics, further emphasizing the effectiveness of ESC-Eval. More correlation experimental results are in Appendix C.
@@ 3.4 ESC-RANK
To facilitate subsequent research, based on InternLM2-7B-ChatCai et al. (2024) and using the manually annotated data in this article, we trained ESC-RANK. ESC-RANK can score the results of multiple rounds of dialogues of different models to our well-designed dimension.
We randomly divided the annotated data into a training set, validation set, and test set according to 7:1:2. Compared with the base model and GPT-4, the results are shown in Table 4.
@@ 4.1 Emotion Support Conversation
Traditional research Sharma et al. (2020); Medeiros and Bosse (2018); Rashkin et al. (2018) on emotion support systems initially focused on simple single-turn emotion dialogue systems. With the emergence of the ESConv Liu et al. (2021) dataset, the development of ESC shifted towards more complex multi-turn dialogues. Researchers have proposed various optimization strategies on ESConv dataset, Peng et al. (2022) introduced an innovative hierarchical graph network, aiming to effectively utilize both the global emotion cause and the local user intention in emotional support conversations. Moving away from relying on a single strategy for response generation, Tu et al. (2022) incorporated commonsense knowledge and a mix of response strategies into the framework of emotional support conversation and so on. With the development of LLMs, which have naturally become suitable for chatbot scenarios due to their generative architecture. Researchers Zheng et al. (2023b); Qiu et al. (2023); Liu et al. (2023) have utilized these models by pertaining and fine-tuning through supervised learning. For instance, Zheng et al. (2023b) used ChatGPT to generate data for constructing an emotion-supported dialogue system, while Madani et al. (2024) expanded the ESconv dataset to address the issue of extrapolating the length capabilities of large language models. In addition, some studies Hua et al. (2024); Zhang et al. (2024); Chen et al. (2023a) have also used LLMs in ESC-related fields, such as psychological counseling. The purpose of our study is to provide a comprehensive and rigorous evaluation of these LLM-based ESC models.
@@ 4.2 Role Play Agents
Recent advancements in LLMs have significantly boosted the rise of Role-Playing Language Agents (RPLAs) Chen et al. (2024), existing researches Wang et al. (2024b); Tu et al. (2024); Shen et al. (2024); Wang et al. (2024a) have proposed multiple evaluation datasets for role-playing, wherein various approaches Li et al. (2023); Shao et al. (2023); Wang et al. (2024b); Zhou et al. (2023) such as In-Context-Learning (ICL) Min et al. (2022), Chain-of-Thought (CoT) Wei et al. (2022) and Supervised Fine-Tuning (SFT) have been explored to construct role-playing models. Additionally, the industry has witnessed the emergence of numerous role-playing products, like Character AI [^3] https://character.ai/ and Reflection AI [^4] https://reflectionai.xyz/, leading to a wide-ranging impact. RPLAs are capable of assuming specific roles, engaging in human-like interactions through composite character settings, role background knowledge, and speech styles, thereby exhibiting human-like attributes and playing a role in everyday conversational contexts. This paper followed the main idea of evaluating ESC models through RPLAs.

### v2

@@ Abstract
Emotion Support Conversation (ESC) is a crucial application, which aims to reduce human stress, offer emotional guidance, and ultimately enhance human mental and physical well-being. With the advancement of Large Language Models (LLMs), many researchers have employed LLMs as the ESC models. However, the evaluation of these LLM-based ESCs remains uncertain. Inspired by the awesome development of role-playing agents, we propose an ESC Evaluation framework (i.e., ESC-Eval), which uses a role-playing agent to interact with ESC models, followed by a manual evaluation of the interactive dialogues. In detail, we first re-organize 2,801 role-playing cards from seven existing datasets to define the roles of the role-playing agent. Second, we train a specific role-playing model - ESC-Role to mimic the behavior of a real person experiencing distress. Third, through ESC-Role and organized role cards, we systematically conduct experiments using 14 LLMs as the ESC models, including general AI-assistant LLMs (e.g., ChatGPT) and ESC-oriented LLMs (e.g., ExTES-Llama). We conduct comprehensive human annotations on interactive multi-turn dialogues of different ESC models. The results show that ESC-oriented LLMs exhibit superior ESC abilities compared to general AI-assistant LLMs, but there is still a gap behind human performance. Moreover, to automate the evaluation of future ESC models, we developed ESC-RANK, which trained on the annotated data, achieving a scoring performance surpassing 35 points of GPT-4. Our data and code are available at https://github.com/haidequanbu/ESC-Eval.
@@
Haiquan Zhaothanks: Work done during internship at Shanghai Artificial Intellignece Laboratory1,2   Lingyu Li2   Shisong Chen1   Shuqi Kong2    Jiaan Wang 1 Kexin Huang1,2   Tianle Gu2   Yixu Wang1,2   Jian Wang2   Dandan Liang2 Zhixu Lithanks: Correspondence to: Zhixu Li<zhixuli@fudan.edu.cn> and Yan Teng <tengyan@pjlab.org.cn>1  Yan Teng22footnotemark: 22    Yanghua Xiao1   Yingchun Wang2 1 Fudan University 2 Shanghai Artificial Intelligence Laboratory
@@ 2.2 Role Card Acquisition
To ensure the diversity of character cards, drawing inspiration from ESConvLiu et al. (2021), ExTESZheng et al. (2023b), and the Life Events ScaleWethington (2016), we first construct a classification system consisting of three hierarchical layers and encompassing 37 categories. Then we propose reconstructing role cards in open-resources data and identifying each role card within each category. The construction of this procedure involves three primary steps. First, we collect 7 open-source datasets that cover a wide range of potential user roles. Then we utilize GPT-4 to extract roles from these datasets and filter out low-quality role cards, followed by human filtering. Finally, we employ a manual annotation process to ensure the quality of the role cards and classify them into their respective tertiary categories. We will introduce each step in the following, and more details can be found in Appendix A.
@@ 2.2.1 Dataset collection
To obtain a diverse set of character cards, we conduct a comprehensive investigation into existing datasets in the field of emotion support and mental health datasets. Subsequently, we select seven datasets as the source datasets for this study. The open-source datasets utilized are listed in Appendix A.
@@ 2.2.2 User cards extraction and filtering
After obtaining these datasets, we encounter both Multi-turn Dialogue (MD) datasets and single-turn Question-and-Answer (QA) datasets. To extract user profiles from these diverse datasets, we employed different prompts for QA and MD datasets using GPT-4 for the initial extraction. The utilization of GPT-4 in this process incurred approximately a cost of \$120. After acquiring the initial character cards, we employ GPT-4 to conduct an initial filtration process on role cards, eliminating those that solely consist of emotions without any associated events. The utilization of GPT-4 in this process incurred an approximate cost of \$70. The prompt used in this section can be found in Appendix A. After the filtering process of GPT-4, we apply a human filter to ensure the quality of these cards.
@@ 2.2.3 Manual annotation and correction
After obtaining role cards that had undergone preliminary screening by GPT-4 and human filter, we employ a two-stage approach involving crowdsourced annotation followed by manual correction to ensure the quality of the role cards.
Crowd workers annotation In the above section, we develop a comprehensive three-tier classification system comprising a total of 37 categories of real-life questions, which are listed in Table 9. Based on this classification, the crowd workers are instructed to annotate the filtered character cards with their corresponding tertiary classifications. This requirement is motivated by the need for convenient evaluation and the quality of role-playing. Additionally, we request the crowd workers to label the high-quality and medium-quality character cards within the dataset. The annotation rules and classification system for annotation can be found in Appendix A. The estimated duration for the annotation phase required is approximately 10 days with 10 crowdsourcing workers.
Human correction Upon completion of the first-stage crowdsourced annotation, we proceed with a second-stage manual correction. We request authors of this paper who are more familiar with this task to conduct an examination of the annotations for each role card, rectifying any incorrect categorizations and addressing issues pertaining to the quality of the role cards
Following the two-stage process of crowdsourced annotation and manual correction, the role cards representing various real-world individuals with different problems are successfully reconstructed. The data analysis of these role cards is listed in Appendix A.
@@ 2.3 ESC-Role
To construct a more robust role-playing model, we train a specific role-playing agent called ESC-Role using both general data and data specific to ESC scenarios for ESC-Eval. The following sections outline the steps involved in training and evaluating this model.
@@ 2.3.1 Data Collection
Using the same procedure as in Section 2.2, we select Smile, ESConv, and ExTES datasets mentioned previously to collect ESC scenarios data. We employ methods including extraction through GPT-4, filtering with GPT-4, and manual filtering to extract role cards from multi-turn dialogues, Resulting in a total of 3,390 role-playing data which consist of a role card and a corresponding dialogue. The role cards are used as system prompts for model training. To further enhance the model’s robust role-playing ability, we filter five role-playing datasets consisting of multi-turn dialogues from Huggingface [^1] https://huggingface.co/. After processing, we acquire 14K role-playing data instances, consisting of both general role-playing instruction data and ESC role-playing data.
@@ 2.3.2 Implementation and Evaluation Metric
Due to the inclusion of both English and Chinese in the character cards, we select Qwen1.5-14B-Chat as our base model. We adopt LoRAHu et al. (2021) parameter-efficient Finetuning on the dataset collected above. We compare ESC-Role with some state-of-the-art role-playing agents like GPT-4 and BaichuanNPC, these agents are API-based LLMs, and we conduct all kinds of prompts like Chain-of-Thought (CoT)Wei et al. (2022) and In-Context-Learning (ICL)Min et al. (2022), more details can refer to Appendix B.
@@ 2.3.3 Evalution Results
In addition, we select pairs that had the best performance with different role-playing models and the source multi-turn dialogue data of role cards. We manually evaluate which dialogue more closely resembled real human conversations. The results are shown in Figure 3. From the figure, it can be observed that it is difficult for humans to distinguish between the results generated by ESC-Role and the results from the original data. Both of them outperform GPT-4 and Baichuan-NPC, demonstrating the effectiveness of using ESC-Role for role-playing in ESC-Eval.
@@ 3 Evaluation
In this section, we conduct evaluations on 14 general LLMs and domain-specific LLMs on ESC-Eval. We first introduce the models for evaluation. Then we display our experimental results. Finally, we display the details of our scoring model ESC-RANK.
@@ 3.1 Evaluating models
We select 14 models for evaluation, including closed-source, open-source, and domain-specific models, which are as follows:
@@
To facilitate a more accurate comparison of the capabilities of various models, we choose models of similar magnitudes, such as the 6B/7B/8B model parameter sizes for comparison.
@@ 3.2 Evaluation Results
Based on pre-defined dimensions, we conduct a comprehensive manual assessment, and the results are presented in Table 2. Both in English and Chinese ESC conditions, domain-specific LLMs (ChatCounselor and EmoLLM), respectively achieved the best results. From Table 2 above, in the comparison between general models and domain-specific models, the general models perform better in terms of fluency, expression diversity, and emotional comfort skills. This can be attributed to their highly structured output, such as phrases like “I understand you very well, it is very normal to feel …, here are some possible suggestions: ” The general models generate a large amount of text, scoring high in terms of advice effectiveness and expression diversity. Besides, due to Larger scale parameters, the API-based models exhibit greater knowledge of emotional comfort, with GPT-4 and ChatGPT demonstrating the highest proficiency. However, these models perform poorly in terms of human-like and human-centric responses, as users in this context expect replies that are more humanized and possess greater human-like qualities. In the comparison of domain-specific models, MindChat, SoulChat, and EmoLLM, which are not fine-tuned in English, showed inferior fluency. On the other hand, ExTES-llama and ChatCounselor perform well. ExTES is fine-tuned with data generated by ChatGPT, while ChatCounselor is fine-tuned using real psychological counseling data, exhibiting superior performance. From Table 2 bottom, the general models perform well in terms of expressing diversity and providing effective suggestions. Trained on diverse and abundant data, EmoLLM exhibits excellent performance across multiple dimensions among the various domain-specific models. Other domain-specific models, due to their remarkable human-like qualities and human convenience, surpass the general models. However, there is still room for improvement in terms of emotional support knowledge, and significant potential exists for enhancing human convenience. It is worth noting that MindChat, trained on bilingual data, not only demonstrates strong Chinese language proficiency but also exhibits commendable English language capabilities.
@@ 3.3 Correlation Analysis
To validate the effectiveness of ESC-Eval, we randomly select 20 instances from the ESConv dataset. We choose three categories from the target model and included five different models for correlation analysis. These models are subjected to interactions with human evaluators who model seekers looking for help. And they are asked to provide ratings upon completion of the interactions, according to human evaluation methods in other papers. The human-rated scores are considered as the optimal evaluation method, and we conduct a correlation analysis between various automatic evaluation methods and the ESC-Eval method. The results are presented in Table 3.
@@
From Table 3, it can be observed that ESC-Eval exhibits the best correlation with the evaluation metrics, except for the Fluency and Empathy indicators. In terms of Fluency, automated metrics outperform ESC-Eval, we analyze that during manual annotation, human annotators may exhibit some bias towards the fluency of segmented statements generated by a general model which significantly deviates from the ESConv dataset. It is observed that human annotators tend to prefer naturally expressed content, leading to relatively lower manual ratings for the fluency outputs of general models. At the same time, the content generated by the general model is quite different from that of ESconv, and the automation metric is also very low. As a result, there is a strong correlation between automated evaluation metrics and humans. However, in ESC-Eval all models perform well on fluency due to the capability of LLMs, leading to low correlation. A similar phenomenon is observed for the Empathy indicator, where although there is some correlation, it is due to the alignment process that most LLMs undergo, which enables them to display decent comforting abilities and analytical skills. In terms of the overall average metric, ESC-Eval demonstrates the most significant correlation compared to the automated metrics, further emphasizing the effectiveness of ESC-Eval. More correlation experimental results are in Appendix C.
@@ 3.4 ESC-RANK
To facilitate subsequent research, based on InternLM2-7B-ChatCai et al. (2024) and using the manually annotated data in this article, we train ESC-RANK. ESC-RANK can score the results of multiple rounds of dialogues of different models to our well-designed dimension.
We randomly divide the annotated data into a training set, validation set, and test set according to 7:1:2. Compared with the base model and GPT-4, the results are shown in Table 4.
@@ 4.1 Emotion Support Conversation
Traditional research Sharma et al. (2020); Medeiros and Bosse (2018); Rashkin et al. (2018) on emotion support systems initially focus on simple single-turn emotion dialogue systems. With the emergence of the ESConv Liu et al. (2021) dataset, the development of ESC shifted towards more complex multi-turn dialogues. Researchers have proposed various optimization strategies on ESConv dataset, Peng et al. (2022) introduced an innovative hierarchical graph network, aiming to effectively utilize both the global emotion cause and the local user intention in emotional support conversations. Moving away from relying on a single strategy for response generation, Tu et al. (2022) incorporate commonsense knowledge and a mix of response strategies into the framework of emotional support conversation and so on. With the development of LLMs, their generative architecture has naturally made them well-suited for chatbot scenarios.Researchers Zheng et al. (2023b); Qiu et al. (2023); Liu et al. (2023) utilize these models by pertaining and fine-tuning through supervised learning. For instance, Zheng et al. (2023b) use ChatGPT to generate data for constructing an emotion-supported dialogue system, while Madani et al. (2024) expand the ESconv dataset to address the issue of extrapolating the length capabilities of large language models. In addition, some studies Hua et al. (2024); Zhang et al. (2024); Chen et al. (2023a) also use LLMs in ESC-related fields, such as psychological counseling. The purpose of our study is to provide a comprehensive and rigorous evaluation of these LLM-based ESC models.
@@ 4.2 Role Play Agents
Recent advancements in LLMs have significantly boosted the rise of Role-Playing Language Agents (RPLAs) Chen et al. (2024). Existing researches Wang et al. (2024b); Tu et al. (2024); Shen et al. (2024); Wang et al. (2024a) have proposed multiple evaluation datasets for role-playing, wherein various approaches Li et al. (2023); Shao et al. (2023); Wang et al. (2024b); Zhou et al. (2023) such as In-Context-Learning (ICL) Min et al. (2022), Chain-of-Thought (CoT) Wei et al. (2022) and Supervised Fine-Tuning (SFT). Additionally, the industry has witnessed the emergence of numerous role-playing products, like Character AI [^3] https://character.ai/ and Reflection AI [^4] https://reflectionai.xyz/, leading to a wide-ranging impact. RPLAs are capable of assuming specific roles, engaging in human-like interactions through composite character settings, role background knowledge, and speech styles, thereby exhibiting human-like attributes and playing a role in everyday conversational contexts. This paper follows the main idea of evaluating ESC models through RPLAs.